# High-Precision Physics Experiments at Huizhou Large-Scale Scientific Facilities

### v1

@@ V.2 Parity-Violating Nuclear Forces from Neutron Capture Measurement
The weak interaction between hadrons remains an unresolved aspect of the Standard Model of particle physics. Experimentally, the only way to study strangeness-conserving weak interactions between hadrons is to measure parity violation in nucleon-nucleon interactions [149, 150]. Due to the weakness of the interaction compared to the strong and electromagnetic forces, parity-violating signals in nucleon-nucleon processes are extremely small and obscured by strong interactions. To date, N-N parity violation has only been observed in a few reaction channels. High-precision measurements using high-intensity polarized neutrons interacting with hydrogen and light nuclei could reveal clear parity-violating signals, advancing our understanding of hadronic properties and completing the Standard Model’s description of weak interactions between hadrons.
@@
Detecting nonzero parity-violating signals is highly challenging. First, sufficient neutron flux ($>10^{16}$ neutrons) is required to achieve statistical significance. Second, systematic errors must be rigorously controlled to eliminate spurious signals. The key technical challenges for the future experiment include: (a) high-intensity, temporally structured cold neutron beams; (b) real-time neutron polarization techniques; (c) stable, high-precision ${\rm{}^{3}He}$-gas target and multi-wire proportional chamber integrated system for neutron capture process and detection of final proton.
@@ VI.1.1 Neutrinoless double-beta decay
For a long time, different phenomenological nuclear models have produced significantly varying NME values, with discrepancies reaching up to a factor of three or more [169, 170], leading to considerable uncertainty in the theoretical interpretation of experimental results. These discrepancies mainly arise from fundamental differences in the treatment of nuclear forces and the truncation of model spaces used to describe nuclear wave functions, making the reduction of model dependence in NME calculations a major challenge. Recent advancements in ab initio nuclear-structure methods have shown promise in providing consistent NME values with theoretical uncertainties for lighter double-beta decay candidates [171, 172, 173]. However, the uncertainties—dominated by many-body truncations—remain substantial [174]. Promising ways to mitigate this problem include refining the approximations of nuclear models and exploring correlations between the NMEs and experimental observables from low-energy nuclear structure experiments. These observables include the double Gamow-Teller strength [175, 176]—a double spin-isospin flip excitation mode that can be probed through high-energy heavy-ion double-charge-exchange reactions [177]—as well as nucleon transfer reactions [178], double-gamma transitions [179], high-energy heavy-ion collisions [180], and properties of nuclear low-lying states [181]. Therefore, precision measurements of nuclear structure properties, such as energy levels and transition probabilities of candidate nuclei, and of reactions at HIAF, will aid in refining nuclear models used to compute the NMEs for neutrinoless double-beta decay.
@@ VII Vortex Beam and Its Physics
There are two main possible approaches to achieving high-energy vortex particle beams: (1) First, applying phase modulation (such as spiral phase plates or holographic gratings) to endow low-energy electron beams with specific phases and intrinsic OAM, followed by accelerating the vortex electron beam to a high-energy state using particle accelerators; (2) Accelerating electrons to achieve a high-energy state first, then using phase modulation or specific scattering to obtain OAM. Both approaches present significant theoretical and technical challenges. In particular, in the second approach, ensuring coherence in high-energy states is difficult, and the requirements for phase modulation materials increase significantly with energy. Vortex particles exhibit unique dynamics in electromagnetic fields within accelerators. However, it has been demonstrated that the manipulation of relativistic twisted charged particles using standard linear accelerators is promising [230, 231, 232, 233, 234]. In comparison, the first approach seems technically viable. However, methods for effectively manipulating or accelerating vortex particles in modern particle accelerators have not been fully understood. Therefore, a comprehensive study of beam dynamics for vortex particles in accelerators has become an urgent priority.
Atomic physics.—
@@
Particle physics.—
One of the promising applications of vortex particles is their use as an alternative to spin-polarized beams in high-energy collisions. Vortex particles can be utilized to study strong interactions, probe hadron spin, and investigate the parton structure of hadrons [225]. High-energy vortex electrons can serve as probes for studying proton structure, where the angular momentum of vortex electrons is transferred to quarks or gluons, enabling new investigations of proton spin. The study of elastic and inelastic scattering of a lepton on a hadron $e_{tw}p_{tw}\rightarrow e’X$, where both particles carry intrinsic orbital angular momentum, may provide a novel approach to analyzing hadron form factors and open new possibilities for quantum tomography of hadron structure. Furthermore, the investigation of deep inelastic scattering in both inclusive and exclusive regimes using vortex particles may lead to the identification of new structure functions that describe the distribution of partons within a proton and their connection to the standard formalism.
@@ VIII Summary
In $\eta$ meson studies, the proposed Super $\eta$ Factory is expected to revolutionize $\eta$ production, achieving statistical yields approximately two orders of magnitude higher than those currently possible at the Super Tau Charm Factory. This increased production capacity will present unparalleled opportunities for discovering new particles and interactions. Detailed investigations of meson decay patterns will facilitate rigorous examinations of fundamental symmetries, including parity and its violations, potentially illuminating the enduring matter-antimatter asymmetry prevalent in the universe. The facility’s unique capabilities will enable precise tests of the Standard Model and enhance our understanding of non-perturbative quantum chromodynamics (QCD), while also shedding light on anomalies such as the Adler-Bell-Jackiw (ABJ) anomaly, which may inform theories on neutrino mass generation under gravitational influences.

### v2

@@ V.2 Parity-Violating Nuclear Forces from Neutron Capture Measurement
The emergence of parity-violating hadron-hadron interactions within the Standard Model remains unsolved and unverified experimentally. Experimentally, the only way to study strangeness-conserving weak interactions between hadrons is to measure parity violation in nucleon-nucleon interactions [156, 157]. Due to the weakness of the interaction compared to the strong and electromagnetic forces, parity-violating signals in nucleon-nucleon processes are extremely small and obscured by strong interactions. To date, N-N parity violation has only been observed in a few reaction channels. High-precision measurements using high-intensity polarized neutrons interacting with hydrogen and light nuclei could reveal clear parity-violating signals, advancing our understanding of hadronic properties and completing the Standard Model’s description of weak interactions between hadrons.
@@
Detecting nonzero parity-violating signals is highly challenging. First, sufficient neutron flux ($10^{16}$ neutrons) is required to achieve statistical significance. Second, systematic errors must be rigorously controlled to eliminate spurious signals. The key technical challenges for the future experiment include: (a) high-intensity, temporally structured cold neutron beams; (b) real-time neutron polarization techniques; (c) stable, high-precision $\rm{}^{3}He$}-gas target and multi-wire proportional chamber integrated system for neutron capture process and detection of final proton.
@@ VI.1.1 Neutrinoless double-beta decay
For a long time, different phenomenological nuclear models have produced significantly varying NME values, with discrepancies reaching up to a factor of three or more [176, 177], leading to considerable uncertainty in the theoretical interpretation of experimental results. These discrepancies mainly arise from fundamental differences in the treatment of nuclear forces and the truncation of model spaces used to describe nuclear wave functions, making the reduction of model dependence in NME calculations a major challenge. Recent advancements in ab initio nuclear-structure methods have shown promise in providing consistent NME values with theoretical uncertainties for lighter double-beta decay candidates [178, 179, 180]. However, the uncertainties—dominated by many-body truncations—remain substantial [181]. Promising ways to mitigate this problem include refining the approximations of nuclear models and exploring correlations between the NMEs and experimental observables from low-energy nuclear structure experiments. These observables include the double Gamow-Teller strength [182, 183]—a double spin-isospin flip excitation mode that can be probed through high-energy heavy-ion double-charge-exchange reactions [184]—as well as nucleon transfer reactions [185], double-gamma transitions [186], high-energy heavy-ion collisions [187], and properties of nuclear low-lying states [188]. Therefore, precision measurements of nuclear structure properties and reaction rates—such as proton-removal cross sections from ($d$, 3He) reactions [189] and two-neutron transfer reactions like ($p$, tt[190]—at HIAF will contribute to refining the nuclear models used to calculate NMEs for neutrinoless double-beta decay.
@@ VII.1 Accelerator physics
There are two main possible approaches to achieving high-energy vortex particle beams: (1) First, applying phase modulation (such as spiral phase plates or holographic gratings) to endow low-energy electron beams with specific phases and intrinsic OAM, followed by accelerating the vortex electron beam to a high-energy state using particle accelerators; (2) Accelerating electrons to achieve a high-energy state first, then using phase modulation or specific scattering to obtain OAM. Both approaches present significant theoretical and technical challenges. In particular, in the second approach, ensuring coherence in high-energy states is difficult, and the requirements for phase modulation materials increase significantly with energy. Vortex particles exhibit unique dynamics in electromagnetic fields within accelerators. However, it has been demonstrated that the manipulation of relativistic twisted charged particles using standard linear accelerators is promising [239, 240, 241, 242, 243, 244, 245]. In comparison, the first approach seems technically viable. However, methods for effectively manipulating or accelerating vortex particles in modern particle accelerators have not been fully understood. Therefore, a comprehensive study of beam dynamics for vortex particles in accelerators has become an urgent priority.
VII.2 Atomic physics
@@ VII.3 Nuclear physics
VII.4 Particle physics
One of the promising applications of vortex particles is their use as an alternative to spin-polarized beams in high-energy collisions. Vortex particles can be utilized to study strong interactions, probe hadron spin, and investigate the parton structure of hadrons [234]. High-energy vortex electrons can serve as probes for studying proton structure, where the angular momentum of vortex electrons is transferred to quarks or gluons, enabling new investigations of proton spin. The study of elastic and inelastic scattering of a lepton on a hadron $e_{tw}p_{tw}\rightarrow e\textquoteright X$, where both particles carry intrinsic orbital angular momentum, may provide a novel approach to analyzing hadron form factors and open new possibilities for quantum tomography of hadron structure. Furthermore, the investigation of deep inelastic scattering in both inclusive and exclusive regimes using vortex particles may lead to the identification of new structure functions that describe the distribution of partons within a proton and their connection to the standard formalism.
@@ VII.4 Particle physics
Vortex particle beams exhibit broad application prospects in the investigation of high-precision problems in atomic, nuclear, and particle physics. The ability to generate high-energy vortex beams is a prerequisite for such studies and represents a key development direction for high-precision experiments at the Huizhou facility. Vortex particle accelerators are among the effective approaches to address this challenge. The feasibility of producing vortex charged particles in linear accelerators by optimizing the magnetic-immersed cathode and stripper foil configuration has been theoretically demonstrated by D. Karlovets and his collaborators [243, 244]. Based on this theoretical framework, the team is currently collaborating with researchers at the Joint Institute for Nuclear Research (JINR) in Dubna, aiming to develop a vortex electron source at the LINAC-200 accelerator to generate vortex electron beams with energies up to 200 MeV.
Similarly, we are planning to conduct research on the generation of vortex ion beams at Huizhou based on the magnetic immersion method. Preliminary results further suggest that, for generating high-quality vortex ion beams, magnetic-immersed stripping should be performed when the ion energy is close to or exceeds 1 MeV/u. For negative hydrogen beams, this corresponds to the output energy of the radio-frequency quadrupole (RFQ) section in the front end of an ion RF accelerator. By placing a room-temperature solenoid with a magnetic field of about 1 T downstream of the RFQ, in the medium-energy beam transport (MEBT) section, it is possible to produce vortex protons with an orbital angular momentum (OAM) quantum number $|\ell|\geq 1$, while maintaining acceptable control over emittance growth. This scheme is also applicable to the generation of vortex heavy-ion beams, although optimization of the initial charge states of the ions is more critical in such cases.
At the Institute of Modern Physics, existing room-temperature ion linac facilities can be upgraded to experimentally verify the vortex ion beam generation using the magnetic-immersed stripping method. The primary challenge at present lies in unambiguously confirming the generation of vortex ions. The development of appropriate diagnostic techniques, along with matching beamline designs, constitutes the next major step in this research.
@@ VIII Summary
In $\eta$ meson studies, the proposed Super $\eta$ Factory is expected to revolutionize $\eta$ production, achieving statistical yields approximately two orders of magnitude higher than those currently possible at the Super Tau Charm Factory. This increased production capacity will present unparalleled opportunities for discovering new particles and interactions. Detailed investigations of meson decay patterns will facilitate rigorous examinations of fundamental symmetries, including parity and its violations, potentially illuminating the enduring matter-antimatter asymmetry prevalent in the universe. The facility’s unique capabilities will enable precise tests of the Standard Model and enhance our understanding of non-perturbative QCD, while also shedding light on anomalies such as the Adler-Bell-Jackiw (ABJ) anomaly, which may inform theories on neutrino mass generation under gravitational influences.
@@
Acknowledgments
This project is supported by National Natural Science Foundation of China under Grant Nos. 12075326; Guangdong Basic and Applied Basic Research Foundation under Grant No. 2025A1515010669; Natural Science Foundation of Guangzhou under Grant No. 2024A04J6243; Fundamental Research Funds for the Central Universities (23xkjc017) in Sun Yat-sen University; and Innovation Training Program for bachelor students in Sun Yat-sen University. The simulation conducted by the MACE working group benefited greatly from the provision of computing resources by the National Supercomputer Center in Guangzhou.